# Energetically consistent localised APE budgets for local and regional studies of stratified flow energetics

### v1

@@ Abstract
Because it allows a rigorous separation between reversible and irreversible processes, the concept of available potential energy (APE) has become central to the study of turbulent stratified fluids. In ocean modelling, it is fundamental to the parameterisation of meso-scale ocean eddies and of the turbulent mixing of heat and salt. However, how to apply APE theory consistently to local or regional subdomains has been a longstanding source of confusion due to the globally defined Lorenz reference state entering the definition of APE and of buoyancy forces being generally thought to be meaningless in those cases. In practice, this is often remedied by introducing heuristic ‘localised’ forms of APE density depending uniquely on region-specific reference states, possibly diverging significantly from the global Lorenz reference state. In this paper, we argue that this practice is problematic because it cannot consistently describes the turbulent APE cascades associated with the inter-scale energy transfers between the APE produced at large scales — which depends on the global Lorenz reference state — and the APE of smaller scales. To avoid this difficulty, we argue that localised forms of APE density should be defined as the eddy APE component of an exact mean/eddy decomposition of the APE density. The eddy APE density thus defined exhibits a much weaker dependency on the global Lorenz reference state than the mean APE, in agreement with physical intuition, but with a different structure than that of existing heuristic localised APE forms. The results are important, because they establish a rigorous physical basis for linking parameterised energy transfers to observable viscous and diffusive dissipation rates, which is a pivotal goal of energetically consistent ocean models.
@@
Exact mean-to-eddy APE conversion can be negative, unlike its approximation
@@ 1 Introduction
Turbulent stratified flows exhibit a complex interplay between reversible and irreversible processes, each providing crucial insights into the other. Reversible aspects are typically associated with adiabatic stirring, which deforms isopycnal surfaces, increasing their areas and magnifying irreversible effects by enhancing tracer gradients. This process leads to the dissipation of mechanical energy and tracer variances at increasingly smaller scales through molecular and diffusive processes (Eckart, 1948). The concept of available potential energy (APE), originally formulated by Margules (1903) and Lorenz (1955) and adapted to the study of turbulent stratified mixing by Winters et al. (1995), serves as a key tool for distinguishing between reversible and irreversible processes. APE theory posits that the potential energy (PE) of any stratified fluid can be partitioned into a component (the APE) available for reversible conversions with kinetic energy (KE) and a component (the background potential energy, BPE) that is not. In Lorenz (1955) approach, the BPE is defined as the PE of a flattened state of minimum potential energy obtainable from the actual state through an adiabatic re-arrangement of mass, effectively representing the ‘heat’ content of the fluid (Tailleux 2025, in preparation). Consequently, APE theory provides a natural framework for assigning distinct energetic signatures to reversible and irreversible processes. Reversible processes affect the APE of the fluid while leaving the BPE unaffected, whereas irreversible processes entail an energy transfer between the APE and BPE. In most cases, the net transfer occurs from the APE to be BPE; however, the reverse conversion is occasionally possible, as observed in double-diffusive instabilities, e.g., Middleton and Taylor (2020); Middleton et al. (2021); Tailleux (2024). In the local theory of APE, the conversion rate between APE and BPE is generally referred to as the APE dissipation rate, denoted as $\varepsilon_{p}$. Although initially introduced in the context of Boussinesq fluids, the nature of $\varepsilon_{p}$ in the fully compressible Navier-Stokes equations has been discussed by Tailleux (2009, 2013c, 2024). In the global APE framework of Winters et al. (1995), the volume-integrated APE dissipation rate is represented by the term $\Phi_{d}-\Phi_{i}$. Together with the viscous dissipation rate $\varepsilon_{k}$, the sum $\varepsilon_{k}+\varepsilon_{p}$ represents the total dissipation of mechanical energy, defined as the sum of KE and APE.
@@
While recent discussions of APE have primarily been framed within the global APE framework of Winters et al. (1995), the concept’s importance and utility were first recognised by Oakey (1982) and Gargett and Holloway (1984) through the derivation of a local APE budget for a quadratic non-negative APE density. This was achieved by rescaling the budget of temperature variance and linking it to the mechanical energy budget. This result was significant for connecting turbulent mixing to the mechanical energy budget, as it possessed a conversion term with kinetic energy and related the APE dissipation rate $\varepsilon_{p}$ to the dissipation of temperature variance
@@
In (1), $g$ represents the acceleration of gravity, $\alpha$ is the thermal expansion coefficient, $\kappa_{T}$ is the thermal diffusivity, $\theta$ is potential temperature with mean $\overline{\theta}$ and perturbation $\theta^{\prime}$. The relative importance of diffusive and viscous effects in dissipating mechanical energy can be quantified using the dissipation ratio $\Gamma=\varepsilon_{p}/\varepsilon_{k}$, a commonly used measure of mixing efficiency often considered to be close to $0.2$. Both $\varepsilon_{p}$ and $\varepsilon_{k}$ can be used to define the turbulent diapycnal mixing according to the formula
@@
e.g., Lindborg (2006). Over the past three decades, the somewhat ad-hoc approach to APE developed by Oakey (1982) and Gargett and Holloway (1984) has since been superseded by Winters et al. (1995) exact global APE framework and the the exact finite amplitude local APE framework first developed by Andrews (1981) and Holliday and McIntyre (1981), and subsequently developed by Shepherd (1993), Scotti et al. (2006), Roullet and Klein (2009), Tailleux (2013b), Scotti and White (2014), Zemskova et al. (2015), and Tailleux (2018), among others.
Reversible and irreversible effects can also be described in terms of energy transfers between different scales of motion. In the atmosphere, there has been extensive discussion about the physical explanation for the turbulent energy cascade affecting both kinetic and potential energy. Lindborg (2006) has developed a theory suggesting a forward energy cascade with spectra given by
@@
with $C_{1}\approx C_{2}$. Interestingly, this theory predicts that the potential energy to kinetic energy ratio at all scales is
$\frac{E_{P_{h}}}{E_{K_{h}}}\approx\frac{\varepsilon_{p}}{\varepsilon_{K}}=\Gamma$ (4)
and therefore determined by the dissipation ratio $\Gamma$.
Lorenz introduced the global APE framework to explain how the atmospheric circulation is maintained against dissipation. He introduced the Lorenz Energy Cycle (LEC) as a means to achieve this, partitioning the KE and APE reservoirs into mean and eddy components. However, because the integrand defining the APE is not positive definite, Lorenz had to rely on various manipulations and integrations by parts to rewrite the volume integrated APE as the volume integral of a quadratic quantity, that could be then split into mean and eddy components. Setting aside compressible effects, the local APE framework defines the APE density of a fluid parcel in terms of the work against buoyancy forces relative to the reference density profile $\rho_{0}(z)$, characterising the globally defined Lorenz reference state of minimum potential energy. However, this reference state is rarely considered relevant for understanding the energetics of much smaller subdomains. Rather, it is generally implicitly assumed that buoyancy forces should be defined relative to some locally defined averaged density field characteristic of the local environment. To continue using APE theory in such cases, most existing approaches appear to be based on some ad-hoc heuristic ‘localisation’ of the local or global APE frameworks. A typical example of such localisation is Roullet et al. (2014) ad-hoc definition of eddy APE, which physically modifies the exact form of local APE density
$E_{a}^{exact}=\frac{g}{\rho_{\star}}\int_{z_{r}(\rho)}^{z}[\rho-\rho_{0}(\tilde{z})]\,{\rm d}\tilde{z}$ (5)
@@
with $\rho_{m}(x,y,z)$ representing some locally defined mean density field, and $z_{m}$ the level of neutral buoyancy satisfying $\rho_{m}(x,y,z_{m})=\rho$. While plausible and physically appealing, (6) is the source of much confusion in the literature about turbulent stratified mixing, as it no longer provides clarity on how to compute the reference state associated with the exact formula (5). Several studies have discussed the issue, e.g., Arthur et al. (2017); Wykes et al. (2015); Dewar and McWilliams (2019) and found that different choices of reference states often lead to significantly different conclusions about the properties of mixing.
@@
Part of the difficulties or confusion surrounding this issue seems to arise from the insistence on discussing the energetics of individual mixing events independently of the energetics of the global ocean in which they are embedded. However, in a statistically steady state, the total KE+APE dissipation must balance the work done by the surface wind stress plus the APE production by the surface buoyancy fluxes. Considering that the APE production by surface buoyancy fluxes is always based on the exact APE density based on Lorenz reference density profile $\rho_{0}(z)$, it seems evident that if the APE dissipation rate based on (6) is sensitive to the choice of $\rho_{m}$, then there must be only one consistent way to define $\rho_{m}$ that can achieve the desired balance. The question addressed in this paper is: which one it is? To a large extent, a resolution to this problem already exists and consists in defining the eddy APE as part of a mean/eddy decomposition of the local APE density (5), as first proposed by Scotti and White (2014). However, the fact that this approach resolves the issue does not appear to be widely recognised. The main obstacle is likely due to the fact that in the majority of studies, the energetics of turbulent stratified mixing continue to be discussed in terms of the global APE framework of Winters et al. (1995) instead of the local APE density framework. A second obstacle is due to the fact that the budgets derived by Scotti and White (2014) often appear more complicated and less physically intuitive than those of Winters et al. (1995). One key aim of this paper, building upon the previous work by Tailleux (2024), is to demonstrate that it is possible to significantly simplify these derivations, making them much simpler than those of Winters et al. (1995) and Scotti and White (2014). We hope this will help promote the wider adoption of the local APE framework as revisited in the present paper.
@@
This paper is organised as follows. Section 2 highlights the convexity of the local APE density as the fundamental property underlying the construction of the concept of eddy APE in the most general case. The convexity property of the local APE density was only briefly mentioned by Scotti and White (2014) but arguably warrants to be more fully discussed and exploited. This section also clarifies the links between exact and heuristic forms of local APE density. Section 3 revisits and simplifies the derivation of the local budgets of mean and eddy APE previously obtained by Scotti and White (2014). Section 4 discusses the constraints on mixing parameterisations deriving from the consideration of the eddy APE and KE budgets, a key issue for the development of energetically consistent numerical ocean models, which do not appear to have been considered before. Section 5 provides a summary and discussion of the results.
@@ 2.1 Boussinesq model equations
This buoyancy $b_{\ell}$ differs from the standard buoyancy $b_{bou}=-g(\rho-\rho_{\star})/\rho_{\star}$ as it is measured relative to the variable reference density $\rho_{0}(z)$ rather than a constant Boussinesq density $\rho_{\star}$. For the purpose of this analysis, we assume that the overall domain, analogous to oceanic conditions, is sufficiently large that the reference density profile $\rho_{0}(z)$ can be considered time-independent. This assumption is supported by climatological observations of temperature and salinity over a century (not showed), which indicate that below the mixed layer, the Lorenz reference state remains stable over time.

### v2

@@ Abstract
Because it allows a rigorous separation between reversible and irreversible processes, the concept of available potential energy (APE) has become central to the study of turbulent stratified fluids. In ocean modelling, it is fundamental to the parameterisation of meso-scale ocean eddies and of the turbulent mixing of heat and salt. However, how to apply APE theory consistently to local or regional subdomains has been a longstanding source of confusion due to the globally defined Lorenz reference state entering the definition of APE and of buoyancy forces being generally thought to be meaningless in those cases. In practice, this is often remedied by introducing heuristic ‘localised’ forms of APE density depending uniquely on region-specific reference states, possibly diverging significantly from the global Lorenz reference state. In this paper, we argue that across-scale energy transfers can only be consistently described if localised forms of APE density are defined as the eddy APE component of an exact mean/eddy decomposition of the APE density, for which a new physically more intuitive and mathematically simpler framework is proposed. The eddy APE density thus defined exhibits a much weaker dependency on the global Lorenz reference state than the mean APE, in agreement with physical intuition, but with a different structure than that of existing heuristic localised APE forms. Our framework establishes a rigorous physical basis for linking parameterised energy transfers to molecular viscous and diffusive dissipation rates. We illustrate its potential usefulness by discussing the energetics implications of standard advective and diffusive parameterisations of the turbulent density flux, which reveals potential new sources of numerical instability in ocean models.
@@
Exact mean-to-eddy APE conversion can be negative, unlike QG approximation
@@ 1 Introduction
Turbulent stratified flows exhibit a complex interplay between reversible and irreversible processes, each providing crucial insights into the other. Reversible aspects are typically associated with adiabatic stirring, which deforms isopycnal surfaces, increases their areas, and magnifies irreversible effects by enhancing tracer gradients. This process leads to the dissipation of mechanical energy and tracer variances at increasingly smaller scales through molecular and diffusive processes (Eckart, 1948). The concept of available potential energy (APE), originally formulated by Margules (1903) and Lorenz (1955) and later adapted to the study of turbulent stratified mixing by Winters et al. (1995), serves as a key tool for distinguishing between reversible and irreversible processes. APE theory posits that the potential energy (PE) of any stratified fluid can be partitioned into a component (the APE) available for reversible conversions with kinetic energy (KE), and a component (the background potential energy, BPE) that is not. In Lorenz’s approach, the BPE is defined as the PE of a flattened state of minimum potential energy obtainable from the actual state through an adiabatic rearrangement of mass Lorenz (1955). Consequently, APE theory provides a natural framework for assigning distinct energetic signatures to reversible and irreversible processes (Butler et al., 2013). Reversible processes affect the APE of the fluid while leaving the BPE unaffected, whereas irreversible processes entail an energy transfer between the APE and BPE. In most cases, the net transfer occurs from the APE to the BPE; however, the reverse conversion is occasionally possible, as observed in double-diffusive instabilities (e.g., Middleton and Taylor (2020); Middleton et al. (2021); Tailleux (2024)). In the local theory of APE, the conversion rate between APE and BPE is generally referred to as the APE dissipation rate, denoted as $\varepsilon_{p}$. Although initially introduced in the context of Boussinesq fluids, the nature of $\varepsilon_{p}$ in the fully compressible Navier-Stokes equations has been discussed by Tailleux (2009, 2013c, 2024). In the global APE framework of Winters et al. (1995), the volume-integrated APE dissipation rate is represented by the term $\Phi_{d}-\Phi_{i}$. Together with the viscous dissipation rate $\varepsilon_{k}$, the sum $\varepsilon_{k}+\varepsilon_{p}$ represents the total dissipation of mechanical energy, defined as the sum of KE and APE.
@@
While recent discussions of APE have primarily been framed within the global APE framework of Winters et al. (1995), the concept’s importance and utility were first recognised by Oakey (1982) and Gargett and Holloway (1984) through the derivation of a local APE budget for a quadratic, non-negative APE density. This was achieved by rescaling the budget of temperature variance and linking it to the mechanical energy budget. This result was significant for connecting turbulent mixing to the mechanical energy budget, as it introduced a conversion term with kinetic energy and related the APE dissipation rate $\varepsilon_{p}$ to the dissipation of temperature variance:
@@
In (1), $g$ represents the acceleration due to gravity, $\alpha$ is the thermal expansion coefficient, $\kappa_{T}$ is the thermal diffusivity, and $\theta$ is the potential temperature with mean $\overline{\theta}$ and perturbation $\theta^{\prime}$. The relative importance of diffusive and viscous effects in dissipating mechanical energy can be quantified using the dissipation ratio $\Gamma=\varepsilon_{p}/\varepsilon_{k}$, a commonly used measure of mixing efficiency, often considered to be close to $0.2$. Both $\varepsilon_{p}$ and $\varepsilon_{k}$ can be used to define the turbulent diapycnal mixing according to the formula
@@
(e.g., Lindborg (2006)). Over the past three decades, the somewhat ad-hoc approach to APE developed by Oakey (1982) and Gargett and Holloway (1984) has been superseded by the exact global APE framework of Winters et al. (1995) and the exact finite-amplitude local APE framework first developed by Andrews (1981) and Holliday and McIntyre (1981), and subsequently extended by Shepherd (1993), Scotti et al. (2006), Roullet and Klein (2009), Tailleux (2013b), Scotti and White (2014), Zemskova et al. (2015), and Tailleux (2018), among others.
Reversible and irreversible effects can also be described in terms of energy transfers between different scales of motion. In the atmosphere, there has been extensive discussion about the physical explanation for the turbulent energy cascade affecting both kinetic and potential energy. Lindborg (2006) developed a theory suggesting a forward energy cascade with energy spectra given by
@@
with $C_{1}\approx C_{2}$. Interestingly, this theory predicts that the ratio of potential energy to kinetic energy spectra at all scales is
$\frac{E_{P_{h}}}{E_{K_{h}}}\approx\frac{\varepsilon_{p}}{\varepsilon_{K}}=\Gamma,$ (4)
thus suggesting that the constraint imposed by the dissipation ratio $\Gamma$ extends to significantly larger scales than the dissipation scales. Intriguingly, the ratio $G_{A}/G_{K}$ of the APE production rate $G_{A}$ by surface buoyancy fluxes to the wind power input into geostrophic motions $G_{K}$ also appears to be close to the dissipation ratio $\Gamma$. Indeed, based on Zemskova et al. (2015), $G_{A}\approx 0.5\,{\rm TW}$, and $G_{K}\approx 2-3\,{\rm TW}$, yielding a ratio $G_{A}/G_{K}\approx 0.17-0.25$. Strong constraints on the relative importance of the KE and APE energy spectra and energy transfers therefore appear to exist over a wide range of scales. Understanding the nature of such constraints and exploiting them to constrain the parameterisations of subgridscale processes controlling the unresolved KE and APE energy cascades forms the basis for the development of energetically consistent numerical ocean models.
Figure 1: (a) Schematic illustrating the differences between the exact APE density (black) versus heuristic APE density (orange) as area under the curves, for a parcel with density $\rho$ at depth $z$; (b) Same as for left panel but for the instantaneous exact eddy APE density (black) versus heuristic APE density (orange) associated with density anomalies centred around the mean density $\rho_{m}(z)$ at depth $z$. Note that the black and orange areas can both be approximated as $\rho^{\prime}\zeta^{\prime}/2$ in terms of the same density anomaly $\rho^{\prime}$ but different displacement anomaly $\zeta^{\prime}$.
Lorenz introduced the global APE framework to explain how the atmospheric circulation is maintained against dissipation. He introduced the Lorenz Energy Cycle (LEC) as a means to achieve this, partitioning the KE and APE reservoirs into mean and eddy components. However, because the integrand defining the APE is not positive definite, Lorenz had to rely on various manipulations and integrations by parts to rewrite the volume-integrated APE as the volume integral of a positive definite quadratic quantity that could then be split into mean and eddy components. Setting aside compressible effects, the local APE framework defines the APE density of a fluid parcel in terms of the work against buoyancy forces relative to the reference density profile $\rho_{0}(z)$, characterising the globally defined Lorenz reference state of minimum potential energy. However, this reference state is rarely considered relevant for understanding the energetics of much smaller subdomains. Rather, it is generally implicitly assumed that buoyancy forces should be defined relative to some locally defined averaged density field characteristic of the local environment. To continue using APE theory in such cases, most existing approaches appear to be based on some heuristic ‘localisation’ of the local or global APE frameworks. A typical example of such localisation is the definition of eddy APE by Roullet et al. (2014), which physically modifies the exact form of local APE density
$E_{a}^{exact}=\frac{g}{\rho_{\star}}\int_{z_{0}(\rho)}^{z}[\rho-\rho_{0}(\tilde{z})]\,{\rm d}\tilde{z}$ (5)
@@
with $\rho_{m}(x,y,z)$ representing some locally defined mean density field, and $z_{m}$ the level of neutral buoyancy satisfying $\rho_{m}(x,y,z_{m})=\rho$, while $z_{0}(\rho)$ is such that $\rho_{0}(z_{0}(\rho))=\rho$. While plausible and physically appealing, (6) is the source of much confusion in the literature about turbulent stratified mixing, as it no longer provides clarity on how to compute the reference state associated with the exact formula (5). Several studies have discussed the issue (e.g., Arthur et al. (2017); Wykes et al. (2015); Dewar and McWilliams (2019)) and found that different choices of reference states often lead to significantly different conclusions about the properties of mixing. To help visualise the nature of the problem, the left panel of Fig. 1 illustrates the differences between the exact (black) and heuristic (orange) APE densities predicted by (5) and (6) in a particular example for which $\rho_{m}$ and $\rho_{0}$ differ from each other, as is normally the case. The right panel anticipates the results of this paper, comparing the exact form of instantaneous eddy APE density resulting from an exact mean/eddy decomposition (black) with the heuristic APE density (orange, same as in left panel), associated with fluctuations around the mean density $\rho_{m}$. This schematic shows that while both approaches are associated with the same density anomalies $\rho^{\prime}$, they are associated with different displacement anomalies $\zeta^{\prime}$, potentially leading to significant differences.
@@
Part of the difficulty or confusion surrounding this issue seems to arise from the insistence on discussing the energetics of individual mixing events independently of the energetics of the global ocean in which they are embedded. However, in a statistically steady state, the total KE+APE dissipation must balance the work done by the surface wind stress plus the APE production by the surface buoyancy fluxes. Considering that the APE production by surface buoyancy fluxes is always based on the exact APE density using the Lorenz reference density profile $\rho_{0}(z)$, it seems evident that if the APE dissipation rate based on (6) is sensitive to the choice of $\rho_{m}$, then there must be only one consistent way to define $\rho_{m}$ that can achieve the desired balance. The question addressed in this paper is: which one is it?
To ensure consistency between APE production at large scales and APE dissipation at molecular scales, the solution discussed in this paper is to define localised eddy APE densities in terms of the eddy component of an exact mean/eddy decomposition of the local APE density. As shown in this paper, this is especially important and necessary where the actual state departs significantly from the Lorenz reference state and hence where the standard QG approximation becomes inaccurate. So far, a general theory for how to do so has been lacking, as existing results have been limited to a Boussinesq fluid with a linear equation of state (Scotti and White, 2014) or a dry atmosphere (Novak and Tailleux, 2018), using methods too specialised to be easily generalisable. It follows that to apply existing theory to a realistic ocean (e.g., Zemskova et al. (2015) or MacCready and Giddings (2016)), several approximations need to be made to circumvent the difficulties arising from the two-component nature of the equation of state and its thermobaric nonlinearity, thus diminishing its benefits. Perhaps because of this, most studies of the Lorenz energy cycle — while often acknowledging the existence of Scotti and White (2014) — still continue to rely on the standard QG approximation used by von Storch et al. (2012) and others.
In this paper, we revisit Scotti and White (2014)’s approach by making it more physically intuitive, mathematically simpler, and more easily generalisable to more complex fluids. To that end, we adopt as our starting point the local budget of APE density developed by Tailleux (2024), which more naturally generalises to two-component fluids. To summarise, our approach defines the mean APE density $E_{a}(\overline{\rho},z)$ as the APE density of the mean density field $\overline{\rho}$, the eddy APE density as the residual $E_{a}^{t}=\overline{E_{a}}-E_{a}(\overline{\rho},z)$, and proves the positive definite character of $E_{a}^{t}$ by linking it to the convexity of $E_{a}(\rho,z)$ with respect to $\rho$. The generality of such results suggests that for a realistic thermobaric ocean, the mean APE density should be defined as $E_{a}^{m}=E_{a}(\overline{S},\overline{\theta},z)$ and hence the eddy APE density as the residual $E_{a}^{t}=\overline{E_{a}}-E_{a}(\overline{S},\overline{\theta},z)$, which is quite different from Zemskova et al. (2015) or MacCready and Giddings (2016), thus paving the way for improving the description of the Lorenz energy cycle in ocean models, as will be developed and discussed in subsequent work.
By interpreting the mean and eddy fields as the resolved and unresolved components of a numerical ocean model, we use our framework to discuss and understand the energetic implications of different possible parameterisations of the turbulent fluxes affecting the across-scale KE and APE energy transfers, similarly to the energetically consistent modelling approach of Carsten Eden (Eden et al., 2014; Eden, 2015, 2016) (note, though, that in the latter studies, potential energy is partitioned into dynamic and potential enthalpy rather than into APE and BPE). Our main focus is on advective and diffusive parameterisations of the turbulent density flux, which in the oceanic case pertain to meso-scale and small scales, but which in our framework cannot be distinguished. Our discussion reveals that some conclusions about the impact of turbulent parameterisations depend on whether the exact or QG approximation is used; it also suggests that such parameterisations can occasionally convert unresolved APE into resolved APE, potentially revealing a new type of numerical diffusive instability. We also find that if the energy transfers associated with the APE cascade are assumed to scale proportionally to the KE cascade, aspects of the parameterisations of vertical momentum transfer discussed by Greatbatch and Lamb (1990) can be recovered.
In line with classical discussions of the Lorenz energy cycle, our exact mean/eddy decomposition framework relies on conventional Eulerian Reynolds averaging. However, we acknowledge that there is increasing interest in understanding the energetics for the thickness-weighted averaged equations in isentropic or isopycnic coordinates (e.g., Bleck (1985); Aiki et al. (2016); Loose et al. (2023)). How to extend our results to this case is left to future work.
@@
This paper is organised as follows. Section 2 highlights the convexity of the local APE density as the fundamental property underlying the construction of the concept of eddy APE in the most general case. The convexity property of the local APE density was only briefly mentioned by Scotti and White (2014) but arguably warrants a more thorough discussion and exploitation. This section also clarifies the links between exact and heuristic forms of local APE density. Section 3 revisits and simplifies the derivation of the local budgets of mean and eddy APE previously obtained by Scotti and White (2014). Section 4 discusses the constraints on mixing parameterisations derived from the consideration of the eddy APE and KE budgets, a key issue for the development of energetically consistent numerical ocean models, which do not appear to have been considered before. Section 5 provides a summary and discussion of the results.
@@ 2.1 Boussinesq model equations
This buoyancy $b_{\ell}$ differs from the standard buoyancy $b_{bou}=-g(\rho-\rho_{\star})/\rho_{\star}$ as it is measured relative to the variable reference density $\rho_{0}(z)$ rather than a constant Boussinesq density $\rho_{\star}$. For the purpose of this analysis, we assume that the overall domain, analogous to oceanic conditions, is sufficiently large that the reference density profile $\rho_{0}(z)$ can be considered time-independent. This assumption is supported by climatological observations of temperature and salinity over a century (not shown), which indicate that below the mixed layer, the Lorenz reference state appears to remain stable over time.